# Approximation in Hölder spaces

### 03 Jan 2024 v1

@@ Abstract.
For a modulus of continuity $\omega$ and Banach spaces $X,Y$ we introduce and study the subspaces $\dot{\operatorname{VC}}^{0,\omega}_{\Upsilon}(X,Y)$ of vanishing scales $\Upsilon\in\{\operatorname{small},\operatorname{large},\operatorname{far}\}$ of the homogeneous Hölder space $\dot{C}^{0,\omega}(X,Y).$ For a wide class of couples $X$ and $Y$, we characterize the subspaces of functions approximable by smooth and Lipschitz and boundedly supported functions in terms of these three vanishing scales. In the particular case $X=\mathbb{R}^{n},$ we identify the spaces $\dot{\operatorname{VC}}^{0,\omega}_{\Upsilon}(\mathbb{R}^{n},Y)$ with the corresponding vanishing mean oscillation spaces $\mathrm{VMO}^{\omega}_{\Upsilon}(\mathbb{R}^{n},Y)$, thus providing a proof for the density of test functions also on these spaces.
Key words and phrases:
smooth approximation, vanishing mean oscillation, Hölder spaces, Banach spaces
2020 Mathematics Subject Classification:
46T20, 41A29, 42B35, 46B20, 26B35
@@ 1. Introduction and main results
In this article we consider Banach spaces $X=(X,\|\cdot\|_{X})$ and $Y=(Y,\|\cdot\|_{Y})$, moduli of continuity $\omega:(0,\infty)\to(0,\infty)$ and study the following subspaces of uniformly continuous functions.
Definition 1.1.
The homogeneous Hölder space $\dot{C}^{0,\omega}(X,Y),$ associated to the modulus $\omega,$ consists of those functions $f:X\to Y$ such that
@@
Because $Y$ is a Banach space, it is easy to see that $\dot{C}^{0,\omega}(X,Y)$ also becomes a Banach space when equipped with the norm
@@
Dropping the term $\|f(0)\|_{Y}$, the seminorm $\|\cdot\|_{\dot{C}^{0,\omega}(X,Y)}$ becomes a norm when the functions $f\in\dot{C}^{0,\omega}(X,Y)$ are defined modulo constant equivalence classes. In the particular case $\omega(t)=t^{\alpha},$ $\alpha\in(0,1],$ we recover the Hölder spaces $\dot{C}^{0,\alpha}(X,Y).$
We notate
@@
and define the vanishing scales
@@
Then, we define the following subclasses of the homogeneous space $\dot{C}^{0,\omega}(X,Y).$
@@
For each scale $\Upsilon\in\{\operatorname{small},\operatorname{large},\operatorname{far}\},$ we define
$\dot{\operatorname{VC}}^{0,\omega}_{\Upsilon}(X,Y)=\operatorname{V}^{0,\omega}_{\Upsilon}(X,Y)\cap\dot{C}^{0,\omega}(X,Y),$
and their intersection
$\dot{\operatorname{VC}}^{0,\omega}(X,Y):=\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)\cap\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{far}}(X,Y)\cap\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(X,Y).$
Whenever no confusion is possible, we stop labeling the spaces in the norms and further simplify notation with $\|x\|_{X}=\|x\|$ and $\|f(x)\|_{Y}=|f(x)|;$ we also drop the symbols $X,Y$ and write $\dot{C}^{0,\omega}(X,Y)=\dot{C}^{0,\omega}(X)=\dot{C}^{0,\omega}.$
@@
Concerning the above definitions, let us observe the following.
  • For each scale $\Upsilon\in\{\operatorname{small},\operatorname{large},\operatorname{far}\},$ the set $\dot{\operatorname{VC}}^{0,\omega}_{\Upsilon}$ is closed with respect to the $\dot{C}^{0,\omega}$-seminorm. In particular, $(\dot{\operatorname{VC}}^{0,\omega}_{\Upsilon},\|\cdot\|_{\dot{C}^{0,\omega}})$ are Banach spaces of modulo constant equivalence classes.
  • By Lemma 2.1 at the beginning of Section 2 it follows that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{far}}\subset\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}.$ The converse inclusion is not true, as shown by elementary examples. It follows a posteriori that the natural scale $\Upsilon=\operatorname{large}$ becomes artificial for some of our results, see e.g. Theorem 1.5 below. Moreover, in general $\dot{\operatorname{VC}}^{0,\omega}_{\Upsilon}\not=\dot{\operatorname{V}}^{0,\omega}_{\Upsilon}$ and $\dot{\operatorname{V}}^{0,\omega}_{\Upsilon}\not=\dot{\operatorname{V}}^{0,\omega}_{\Upsilon^{\prime}},$ for $\Upsilon\not=\Upsilon^{\prime}.$ Interestingly then, when we take the intersection of all the scales, we have
    $\dot{\operatorname{VC}}^{0,\omega}=\dot{\operatorname{V}}^{0,\omega}:=\dot{\operatorname{V}}^{0,\omega}_{\operatorname{small}}\cap\dot{\operatorname{V}}^{0,\omega}_{\operatorname{far}}\cap\dot{\operatorname{V}}^{0,\omega}_{\operatorname{large}},$
    and for the inclusion $\dot{\operatorname{V}}^{0,\omega}\subset\dot{C}^{0,\omega}$ above, we need the large scale on the right-hand side.
We assume the modulus to be non-decreasing, satisfy the following coercivity properties
@@
and to be doubling: there exists a doubling constant $C_{\operatorname{db}}>1$ such that
@@
We mention explicitly whenever any of the modulus’ properties are used and we state each result with the minimal assumptions used.
For the Hölder moduli $\omega(t)=t^{\alpha},$ the classes $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}$ are sometimes called the little Hölder spaces. These spaces are fundamental in the study of Lipschitz algebras in metric spaces, and we refer to the monograph by Weaver [Weaverbook, Chapters 4 and 8] for background. See also the recent monograph by D. Mitrea, I. Mitrea, and M. Mitrea [Mitreabook, Chapter 3] for a detailed exposition of the Hölder spaces in the setting of Ahlfors regular sets, their connection with functions of vanishing mean oscillation, and for an introduction to the little Hölder spaces.
@@ 1.1. Main results
We establish various theorems on smooth and Lipschitz approximation for functions of the classes $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and $\dot{\operatorname{VC}}^{0,\omega}(X,Y),$ one of the main features being the fact that our approximations are with respect to the $\dot{C}^{0,\omega}(X,Y)$ seminorm. In the last decades, several fundamental results have been established in the theory of uniform approximation of continuous functions by smooth functions. The optimal order of smoothness one can arrange for those approximations is often determined by the regularity of a canonical or an equivalent renorming of the spaces $X$ and $Y.$ Regarding background references on smoothness and renorming on Banach spaces, and the related approximation results, we refer to the monographs [BenyaminiLindenstrauss] by Benyamini and Lindenstrauss; [DevilleGodefroyZizler] by Deville, Godefroy, and Zizler; and [HajJoh10] by Hájek and Johanis. Closely related to our resuls are the articles by Azagra, Fry, and Keener [AzFrKe12], Cepedello-Boiso [Cepedello], Fry [Fry07], Hájek and Johanis [HajJoh09, HajJoh10], Lasry and Lions [LasryLions], Moulis [Moulis] and Pisier [Pisier]. We next clarify that uniform convergence and $\dot{C}^{0,\omega}$ convergence do not follow from each other.
@@
Endow $C^{0,\omega}(X,Y)$ with the complete norm defined in (1). The following claims can be easily verified either directly from the definition of $|||\cdot|||_{C^{0,\omega}}$ or elementary one-dimensional examples.
@@
Morally our main result is the recognition of the scales $\dot{\operatorname{VC}}^{0,\omega}_{\Upsilon}$ as precisely the correct conceptual tool that allows a complete description of the closure of smooth, Lipschitz and/or boundedly supported functions with respect to the $\dot{C}^{0,\omega}$ seminorm. In order for the reader to familiarize with the conclusions of our theorems, we begin by stating a result (without smoothness) on approximation of $\dot{\operatorname{VC}}^{0,\omega}(X,Y)$ mappings by $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ functions with bounded support.
@@
$\dot{\operatorname{VC}}^{0,\omega}(X,Y)=\overline{\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)\cap C_{\operatorname{bs}}(X,Y)}^{[\dot{C}^{0,\omega}(X,Y)]}.$
Here and below, for any class of functions $\mathcal{C}(X,Y)\subset\dot{C}^{0,\omega}(X,Y),$ we denote by $\overline{\mathcal{C}(X,Y)}^{[\dot{C}^{0,\omega}(X,Y)]}$ the closure of $\mathcal{C}(X,Y)$ under the $\dot{C}^{0,\omega}(X,Y)$ seminorm. Naturally, here we refer to the family of those $f\in\dot{C}^{0,\omega}(X,Y)$ for which there exists a sequence $\{f_{n}\}_{n}\subset\mathcal{C}(X,Y)$ so that $\lim_{n}\|f_{n}-f\|_{\dot{C}^{0,\omega}(X,Y)}=0.$ Furthermore, the subscripted set $\mathcal{C}_{\operatorname{bs}}(X,Y)$ consists of those functions $h:X\to Y$ of class $\mathcal{C}$ that have bounded support. In Theorem the class 1.5 $C_{\operatorname{bs}}(X,Y)$ denotes the space of continuous functions $f:X\to Y$ with bounded support, in particular.
While Theorem 1.5 does not provide any order of smoothness in the approximations, it turns out to be a crucial step for the rest of our approximation theorems. For instance, letting $X=\mathbb{R}^{n}$ and $Y$ be an arbitrary Banach space, we combine Theorem 1.5 with elementary mollification arguments to obtain $C^{\infty}_{\operatorname{bs}}(\mathbb{R}^{n},Y)$ approximations as in Theorem 1.10 below. These two theorems will be proved in Section 2.
With this being said, our results are not confined to functions with a finite-dimensional source $X.$ In particular, letting $X$ be a normed space and $Y=\mathbb{R},$ we prove smooth and Lipschitz approximation for the classes $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,\mathbb{R})$ and $\dot{\operatorname{VC}}^{0,\omega}(X,\mathbb{R}).$ (The corresponding approximations for $\mathbb{C}$-valued functions, in place of $\mathbb{R}$-valued, follow on each occassion after splitting them into their real and imaginary parts.) In particular, with $X$ being a separable normed space, we have the following general result.
@@
Let $k\in\mathbb{N}\cup\{\infty\}$ and let $X$ be a separable normed space admitting a $C^{k}$ and Lipschitz bump function. Then, for a modulus $\omega$ satisfying (2), (3) and (4), the following hold:
@@
Notice that, together with the $C^{k}$ order of smoothness, Theorem 1.6 guarantees Lipschitz approximations. Any separable normed space $X$ with an equivalent norm of class $C^{k}$ satisfies the assumption of Theorem 1.6; see Remark 3.6. In the particular case where $X$ is a separable Hilbert space, one can arrange $C^{\infty}$ smooth approximations, thus obtaining an infinite dimensional version of Theorem 1.10; see Corollary 3.7. We will restate Theorem 1.6 and prove it in Subsection 3.1.
Another class of spaces $X$ on which we obtain smooth approximations are super-reflexive Banach spaces. As a consequence of a renorming theorem due to Pisier [Pisier], any super-reflexive space admits a renorming with modulus of smoothness of power type $1+\alpha,$ for some $\alpha\in(0,1].$ See the property (38) below for the precise formulation. This class of spaces includes the Hilbert (separable or not) spaces, together with many of the classical function Banach spaces, such as the $L^{p}$ spaces, with $1<p<\infty;$ see Remark 3.11. The following theorem establishes $C^{1,\alpha}$ and Lipschitz approximations for functions of the classes $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ and $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X).$
@@
Let $X$ be a super-reflexive space, and $\alpha\in(0,1]$ so that $X$ admits a renorming with modulus of smoothness of power type $1+\alpha.$ Then, for a modulus $\omega$ satisfying (2), (3) and (4), the following hold:
@@
It is worth mentioning that during the proof of Theorem 1.7, we have obtained an interesting result concerning uniform approximation by Lipschitz and $C^{1,\alpha}$ functions in super-reflexive spaces. We remind that in the particular case where $X$ is a Hilbert space, the method of Lasry and Lions [LasryLions] permits to obtain uniform approximation of Lipschitz functions by $C^{1,1}$ and Lipschitz functions. For super-reflexive spaces $X$ that are not necessarily Hilbert spaces, M. Cepedello-Boiso [Cepedello] proved, by an ingenuous variant of the Lasry-Lions method, that Lipschitz functions can be uniformly approximated by $C^{1}$ functions whose derivatives are $\alpha$-Hölder continuous on bounded sets, and so these approximations are Lipschitz on bounded sets. In Theorem 1.8 below, we have obtained uniform approximations that are globally $C^{1,\alpha}$ and globally Lipschitz on $X$, preserving the Lipschitz constants up to a multiplicative factor depending only on $\alpha$ and $X.$
@@
Let $X$ be a super-reflexive space with modulus of smoothness of power type $1+\alpha,$ for $\alpha\in(0,1].$ Then, there exists a constant $\kappa=\kappa(\alpha,X)>0$ for which the following holds. Given an $L$-Lipschitz function $f:X\to\mathbb{R}$ and $\varepsilon>0,$ there exists $g\in C^{1,\alpha}(X)$ so that $g$ is $\kappa L$-Lipschitz on $X,$ and $\sup_{X}|f-g|\leq\varepsilon.$
In the proof of Theorem 1.8, we use a recent result on $C^{1,\omega}$ extensions for jets [AzaMud21] to construct $C^{1,\alpha}$ separating functions with several special properties. This is combined with a method of Hájek and Johanis [HajJoh10, Proposition 1] to glue up a suitable sequence of these separating functions. Both Theorems 1.7 and 1.8 will be restated and proved in Subsection 3.2.
Finally, for mappings $f:X\to Y,$ where $Y$ is an arbitrary Banach space, in Subsection 3.3 we obtain $C^{\infty}$ smooth approximations in the case where $X$ is a space of the form $X=c_{0}(\Gamma).$
Theorem 1.9.
For an arbitrary set of indices $\Gamma,$ let $X=c_{0}(\Gamma),$ and let $Y$ be a Banach space. Then, for any modulus $\omega$ satisfying (2), (3) and (4), the following hold:
@@
This theorem will be restated and proved in Subsection 3.3; see Theorem 3.12. There, we will also show that the approximations can be taken to be Lipschitz in the case where $Y=\mathbb{R}.$
1.2. Applications on the Euclidean space
Let $b\in L^{1}_{\operatorname*{loc}}(\mathbb{R}^{n})$ and $T$ be a non-degenerate Calderón-Zygmund operator. Then, with emphasis on recognizing the correct conditions for different arrangements of the eponents $1<p,q<\infty,$ the following picture
[…]
fully characterizes the boundedness of the commutator $[b,T]f:=bTf-T(bf).$ The stated references are Coifman, Rochberg, Weiss [CRW], Janson [Jan1978] and we especially refer to Hytönen [HyLpLq] for an in depth discussion of the above results and their history. The corresponding characterization of the compactness of commutators can be stated as follows. Let $1<p,q<\infty$ and $Z^{p,q}(\mathbb{R}^{n})$ denote the function space read from the right-hand side of the line (5): then, $[b,T]\in\mathcal{K}(L^{p}(\mathbb{R}^{n}),L^{q}(\mathbb{R}^{n}))$ is a compact operator between the given spaces, if and only if $b\in\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{Z^{p,q}(\mathbb{R}^{n})}.$ The case $p=q,$ is due to Uchiyama [Uch1978]; the case $q>p,$ due to Guo, He, Wu and Yang [GHWY21]; and the case $q<p,$ due to Hytönen, Li, Tao and Yang [HLTY2023]. Now we move to further discuss the case $q>p$ that most closely connects with our work.
A particular (easy) corollary of our Theorem 1.5 is the following Theorem 1.10, which we prove in Section 2.

### 21 Feb 2024 v2

@@ Abstract.
For a modulus of continuity $\omega$ and normed spaces $X,Y$ we introduce and study the subspaces $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y)$ of vanishing scales $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\}$ of the homogeneous Hölder space $\dot{C}^{0,\omega}(X,Y).$ For many couples $X$ and $Y$, we characterize the subspaces that admit approximations by smooth, Lipschitz and boundedly supported functions in terms of these three vanishing scales. In the particular case $X=\mathbb{R}^{n},$ we identify $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)=\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y)$ with the corresponding vanishing mean oscillation spaces, thus transferring these approximations also to these spaces.
Key words: smooth approximation, vanishing mean oscillation, Hölder spaces, Banach spaces
2020 Mathematics Subject Classification: 46T20, 41A29, 42B35, 46B20, 26B35
@@ 1. Introduction and main results
We consider normed spaces $X=(X,\|\cdot\|_{X}),$ $Y=(Y,\|\cdot\|_{Y})$ and moduli of continuity $\omega:(0,\infty)\to(0,\infty)$ and study the Hölder spaces $\dot{C}^{0,\omega}(X,Y)$ that consist of those functions $f$ for which
@@
These spaces are subspaces of uniformly continuous functions, in particular. We establish various full characterizations of approximability in the class $\dot{C}^{0,\omega}(X,Y)$ by Lipschitz and/or smooth and/or boundedly supported test functions. This is done in terms of the three vanishing scales that we define below, but before that let us make a few observations. It is easy to see that, if $Y$ a Banach space, then $\dot{C}^{0,\omega}(X,Y)$ also becomes a Banach space when equipped with the norm
@@
By dropping the term $\|f(0)\|_{Y},$ the seminorm $\|\cdot\|_{\dot{C}^{0,\omega}(X,Y)}$ becomes a norm when the functions $f\in\dot{C}^{0,\omega}(X,Y)$ are defined modulo constant equivalence classes. Moreover, in the particular case $\omega(t)=t^{\alpha},$ $\alpha\in(0,1),$ we recover the Hölder spaces $\dot{C}^{0,\alpha}(X,Y).$ Now, let us notate
@@
and then we define the three vanishing scales
@@
Then, we define the vanishing subspaces of $\dot{C}^{0,\omega}(X,Y).$
@@
For each scale $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\},$ let
$\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y)=\dot{\operatorname{V}}^{0,\omega}_{\Gamma}(X,Y)\cap\dot{C}^{0,\omega}(X,Y),$
and
$\dot{\operatorname{VC}}^{0,\omega}(X,Y)=\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)\cap\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{far}}(X,Y)\cap\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(X,Y).$
Whenever no confusion is possible, we drop the source and target spaces in the norms and simplify notation with $\|x\|_{X}=\|x\|$ and $\|f(x)\|_{Y}=|f(x)|;$ we also write $\dot{C}^{0,\omega}(X,Y)=\dot{C}^{0,\omega}(X)=\dot{C}^{0,\omega}.$
@@
A few remarks concerning the above definitions.
  • For $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\},$ the set $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}$ is closed with respect to the $\dot{C}^{0,\omega}$-seminorm. In particular, $(\dot{\operatorname{VC}}^{0,\omega}_{\Gamma},\|\cdot\|_{\dot{C}^{0,\omega}})$ are Banach spaces of modulo constant equivalence classes, provided the target space $Y$ is a Banach space.
  • By Lemma 2.1 at the beginning of Section 2 it follows that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{far}}\subset\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}.$ The converse inclusion is not true, as shown by elementary examples. It follows a posteriori that the natural scale $\Gamma=\operatorname{large}$ becomes artificial for some of our results, see e.g. Theorem 1.4 below. Moreover, in general $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}\not=\dot{\operatorname{V}}^{0,\omega}_{\Gamma}$ and $\dot{\operatorname{V}}^{0,\omega}_{\Gamma}\not=\dot{\operatorname{V}}^{0,\omega}_{\Gamma^{\prime}},$ for $\Gamma\not=\Gamma^{\prime}.$ Interestingly then, when we take the intersection of all the scales, we have $\dot{\operatorname{VC}}^{0,\omega}=\dot{\operatorname{V}}^{0,\omega}_{\operatorname{small}}\cap\dot{\operatorname{V}}^{0,\omega}_{\operatorname{far}}\cap\dot{\operatorname{V}}^{0,\omega}_{\operatorname{large}},$ and for the last inclusion ``""```$\supset$"` ` ⊃ ", we need the large scale on the right-hand side. These facts are not crucial to our results and we leave their verification to the interested reader.
Our modulus satisfies the following properties
@@
and there exists a doubling constant $C_{\operatorname{db}}>1$ such that
@@
We mention explicitly whenever any of the modulus’ properties $(2),(3),(4)$ are used. Moreover, we always assume the modulus $\omega$ to be non-decreasing.
For the Hölder moduli $\alpha(t):=t^{\alpha},$ the classes $\dot{\operatorname{VC}}^{0,\alpha}_{\operatorname{small}}$ are sometimes called the little Hölder spaces. These are fundamental in the study of Lipschitz algebras in metric spaces, and we refer to the monograph by Weaver [Weaverbook, Chapters 4 and 8] for background. See also the recent monograph by D. Mitrea, I. Mitrea, and M. Mitrea [Mitreabook, Chapter 3] for a detailed exposition of the Hölder spaces in the setting of Ahlfors regular sets, their connection with functions of vanishing mean oscillation, and for an introduction to the little Hölder spaces.
@@ 1.1. Main results
During the last half a century, several fundamental results have been established in the theory of uniform approximation of continuous functions by smooth functions. The optimal order of smoothness arrangeable for these approximations is often determined by the regularity of a canonical or an equivalent renorming of the spaces $X$ and $Y.$ For background references on smoothness and renorming on Banach spaces, and the related approximation results, we refer to the monographs [BenyaminiLindenstrauss] by Benyamini and Lindenstrauss; [DevilleGodefroyZizler] by Deville, Godefroy, and Zizler; and [HajJoh10] by Hájek and Johanis. Closely related to our resuls are the articles by Azagra, Fry, and Keener [AzFrKe12], Cepedello-Boiso [Cepedello], Fry [Fry07], Hájek and Johanis [HajJoh09, HajJoh10], Lasry and Lions [LasryLions], Moulis [Moulis] and Pisier [Pisier]. We next clarify that uniform convergence and $\dot{C}^{0,\omega}$ convergence do not follow from each other.
@@
Endow $C^{0,\omega}(X,Y)$ with the norm $|||\cdot|||_{C^{0,\omega}}$ defined in (1). The next claims can be verified either directly from the definition or by elementary one-dimensional examples.
@@
Our main result is the recognition of the scales $\{\operatorname{small},\operatorname{large},\operatorname{far}\}$ as precisely the concept that allows a complete description of the closure of smooth, Lipschitz and/or boundedly supported functions with respect to the $\dot{C}^{0,\omega}$ seminorm. In order for the reader to familiarize with the conclusions of our theorems, we begin by stating a result (without smoothness) on approximations in $\dot{\operatorname{VC}}^{0,\omega}(X,Y)$ by $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ functions with bounded support.
@@
$\dot{\operatorname{VC}}^{0,\omega}(X,Y)=\overline{\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)\cap C_{\operatorname{bs}}(X,Y)}^{\dot{C}^{0,\omega}(X,Y)}.$
Here and below, for any class of functions $\mathcal{C}(X,Y)\subset\dot{C}^{0,\omega}(X,Y),$ we denote by $\overline{\mathcal{C}(X,Y)}^{\dot{C}^{0,\omega}(X,Y)}$ the closure of $\mathcal{C}(X,Y)$ under the $\dot{C}^{0,\omega}(X,Y)$ seminorm. Naturally, here we refer to the family of those $f\in\dot{C}^{0,\omega}(X,Y)$ for which there exists a sequence $\{f_{n}\}_{n}\subset\mathcal{C}(X,Y)$ so that $\lim_{n}\|f_{n}-f\|_{\dot{C}^{0,\omega}(X,Y)}=0.$ The subscripted set $\mathcal{C}_{\operatorname{bs}}(X,Y)$ consists of those boundedly supported $h:X\to Y$ in the class $\mathcal{C}.$ In Theorem 1.4 the class $C_{\operatorname{bs}}(X,Y)$ denotes the space of continuous functions $f:X\to Y$ with bounded support, in particular. Naturally, the functions in $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ are continuous, and the intersection with $C_{\operatorname{bs}}(X,Y)$ indicates that the approximations have bounded support.
While Theorem 1.4 does not provide any smoothness in the approximations, it turns out to be a crucial step for the rest of our approximation theorems. For instance, letting $X=\mathbb{R}^{n}$ and $Y$ be a Banach space, we combine Theorem 1.4 with elementary mollification arguments to obtain $C^{\infty}_{\operatorname{bs}}(\mathbb{R}^{n},Y)$ approximations in Theorem 1.8 below and both theorems will be proved in Section 2.
Furthermore, our results are not confined to functions with a finite-dimensional source $X,$ and letting $X$ be a normed space and $Y=\mathbb{R},$ we prove smooth and Lipschitz approximation for the classes $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,\mathbb{R})$ and $\dot{\operatorname{VC}}^{0,\omega}(X,\mathbb{R}).$ (The corresponding approximations for $\mathbb{C}^{n}$-valued functions, in place of $\mathbb{R}$-valued, follow on each occassion after splitting them into their $n$-coordinates and then each to their real and imaginary parts.) In particular, with $X$ being a separable normed space, we have the following general result.
@@
Let $k\in\mathbb{N}\cup\{\infty\}$ and $X$ be a separable normed space admitting a $C^{k}$ and Lipschitz bump function. Then, for a modulus $\omega$ satisfying (2), (3) and (4), the following hold:
@@
Notice that, together with the $C^{k}$ order of smoothness, Theorem 1.5 guarantees Lipschitz approximations. Any separable normed space $X$ with an equivalent norm of class $C^{k}$ satisfies the assumption of Theorem 1.5, see Remark 3.6. In the particular case where $X$ is a separable Hilbert space, one can arrange $C^{\infty}$ smooth approximations, thus obtaining an infinite dimensional version of Theorem 1.8, see Corollary 3.7. We will restate Theorem 1.5 and prove it in Subsection 3.1.
Another class of spaces $X$ on which we obtain smooth approximations are super-reflexive Banach spaces. A theorem of Pisier [Pisier] says that any super-reflexive space admits a renorming with modulus of smoothness of power type $1+\alpha,$ for some $\alpha\in(0,1].$ See the property (39) below for the precise formulation. This class of spaces includes the Hilbert (separable or not) spaces, together with many of the classical Banach function spaces, such as the $L^{p}$ spaces, with $1<p<\infty;$ see Remark 3.11. The following theorem establishes $C^{1,\alpha}$ and Lipschitz approximations for functions of the classes $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ and $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X).$
@@
Let $X$ be a super-reflexive space, and $\alpha\in(0,1]$ such that $X$ admits a renorming with modulus of smoothness of power type $1+\alpha.$ Then, for a modulus $\omega$ satisfying (2), (3) and (4), the following hold:
@@
Finally, for mappings $f:X\to Y,$ where $Y$ is an arbitrary Banach space, in Subsection 3.3 we obtain $C^{\infty}$ smooth approximations in the case where $X$ is a space of the form $X=c_{0}(\mathcal{A}).$
@@
For an arbitrary set of indices $\mathcal{A},$ let $X=c_{0}(\mathcal{A}),$ and let $Y$ be a Banach space. Then, for any modulus $\omega$ satisfying (2), (3) and (4), the following hold:
@@
This theorem will be restated and proved in Subsection 3.3, see Theorem 3.12. There, we will also show that the approximations can be taken to be Lipschitz in the case $Y=\mathbb{R}.$
1.2. Applications on $X=\mathbb{R}^{n}$
We provide a brief discussion of how our results relate to compactness of commutators of singular integral operators (SIOs). Compact operators are in particular bounded, and for many non-degenerate bounded SIOs $T$ and $p,q\in(1,\infty)$ there holds that
$\|[b,T]\|_{L^{p}(\mathbb{R}^{n})\to L^{q}(\mathbb{R}^{n})}\eqsim\|b\|_{X^{p,q}(\mathbb{R}^{n})},\quad X^{p,q}(\mathbb{R}^{n})=\begin{cases}\dot{C}^{0,\alpha(p,q)}(\mathbb{R}^{n}),&q>p,\\ \mathrm{BMO}(\mathbb{R}^{n})&q=p,\\ \dot{L}^{r(p,q)}(\mathbb{R}^{n}),&q<p,\end{cases}$ (5)
where $[b,T]f=bTf-T(bf)$ is the commutator and $b:\mathbb{R}^{n}\to\mathbb{C}$ is called the symbol of the commutator. The appearing exponents are defined through $\alpha(p,q)=n(1/p-1/q)$ and $1/q=1/r(p,q)+1/p,$ and $\dot{C}^{0,\alpha(p,q)}(\mathbb{R}^{n})$ is the inhomogeneous Hölder space, $\mathrm{BMO}(\mathbb{R}^{n})$ stands for the space of bounded mean oscillations and $\dot{L}^{r(p,q)}(\mathbb{R}^{n})$ is $L^{r(p,q)}(\mathbb{R}^{n})$ modulo additive constants. Emphasizing the recognition of the correct conditions on the symbol, the bound (5) is attributable to Coifman, Rochberg and Weiss [CRW], when $q=p;$ to Janson [Jan1978], when $q>p;$ and to Hytönen [HyLpLq], when $q<p,$ to where we also refer for a history leading to (5).
If $\alpha(p,q)>1,$ when $q>p,$ then $\dot{C}^{0,\alpha(p,q)}(\mathbb{R}^{n})$ consists of the constant functions and the commutator is bounded iff it is the zero operator iff the symbol is constant. Therefore, regarding compactness, only the range $\alpha(p,q)\leq 1$ is interesting to study. When $\alpha(p,q)=1,$ Guo et al. [GHWY21, Theorem 1.6.] showed that the commutator is compact iff it is the zero operator, again. Now, with $p,q\in(1,\infty),$ and such that if $q>p,$ then $\alpha(p,q)<1,$ for a wide class of non-degenerate SIOs there holds that
$[b,T]\in\mathcal{K}\left(L^{p}(\mathbb{R}^{n}),L^{q}(\mathbb{R}^{n})\right)\Longleftrightarrow b\in Y^{p,q}(\mathbb{R}^{n}):=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{X^{p,q}(\mathbb{R}^{n})}.$ (6)
Emphasizing the recognition of the correct conditions on the symbol, (6) is due to Uchiyama [Uch1978], when $p=q$; due to Guo, He, Wu and Yang [GHWY21], when $q>p;$ and due to Hytönen, Li, Tao and Yang [HLTY2023], when $q<p.$ We remark that it was noted recently [Oik24] that in (6) the sufficiencies of $Y^{p,q}(\mathbb{R}^{n})$ for commutator compactness in all three cases ``$q<p",$`$q=p"$ and ``$q>p"$ all follow easily from each other by classical interpolation of compactness methods. To obtain also the necessity directions on the line (6), it is important to have descriptions of $Y^{p,q}(\mathbb{R}^{n}),$ in the cases $q\geq p,$ in terms of vanishing mean oscillation criteria. These criteria follow immediately from the existence of approximations, but to establish approximations from vanishing mean oscillation criteria takes more work. The existence of nice approximations beginning from vanishing type criteria is exactly the topic of the present article, remarkably. A particular corollary of our Theorem 1.4 is the following theorem.